OpenStudy (anonymous):

[2.06] Solve the equation 2 over 5 (x − 4) = 2x for x. −8 negative 1 over 2 1 −1

OpenStudy (owlcoffee):

\[\frac{ 2 }{ 5(x-4) }=2x\] As you can see, we have a proportion, so, let's first do distributive on that denominator. \[\frac{ 2 }{ 5x-20 }=2x\] So, let's pass the whole (5x-20) multiplying to the other side: \[2=2x(5x-20)\] with that done, let's apply distributive there: \[2=10x ^{2}-40x\] Let's pass that 2 as a negative so we can have everything equal to zero: \[10x ^{2}-40x-2=0\] Let's apply the general formula for this situation: \[x=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\] Where: a=10 b=-40 c=-2 It should look like this: \[x=\frac{ -(-40)\pm \sqrt{(-40)^{2}-4(10)(-2)} }{ 2(10) }\] Think you can take over from here? :)

OpenStudy (anonymous):

5x^2-20x-1=0 x^2-4x-1/5=0 make it square x^2-4x+4-4-1/5=0 (x-2)^2-21/5=0 (x-2)^2=21/5 x-2=+-(sqrt(21/5)) x1=sqrt(21/5) x2=-sqrt(21/5)